# Large deviations for the smallest eigenvalue of a deformed GOE with an outlier

### v1

@@ 2.6. Solving the functional equation in the region $\omega_{\nu,t}(\lambda)\leq\Lambda$
where $\omega^{*}_{\nu,t}(\lambda)$ stands for the unique solution $\gamma$ in $[\omega_{\nu,t}(\ell_{\nu,t}),\inf\eta_{k})$ of the equation $H_{\nu,t}(\gamma)=\lambda$. Also define
@@
and set ${I_{\nu,t}^{\Lambda}(\lambda):=I_{\nu,t}^{\Lambda}(\lambda,\gamma(\Lambda,\lambda))}$. We first prove that $F_{\nu,t}^{+}$ and $F_{\nu,t}^{-}$ equal $I_{\nu,t}^{\Lambda}$ in the simplest case where the functional equations (2.19) and (2.25) are such that the infimum is taken when the term depending on $F_{\nu,t}^{+}$ or $F_{\nu,t}^{-}$ vanishes.
@@
Let $\Lambda\leq\ell_{\nu}$ and $\lambda<\ell_{\nu,t}$. We have for every $\lambda<\ell_{\nu,t}$ and $\Lambda\leq\ell_{\nu}$,
@@
Moreover, if $\Lambda\geq\omega_{\nu,t}(\lambda)$ or $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$, then
@@
A key point to prove Lemma 2.8 is to study the minimizers of $K_{\nu,t}(\Lambda,\lambda,.)$, namely the minimizers of the function $J_{\nu,t}^{\Lambda}(\lambda,.)$ defined in (2.17).
@@
Fix $\lambda\in(-\infty,\ell_{\nu,t}]$.
@@
    $y_{i}(\gamma):=\frac{t}{(\eta_{i}-\omega_{\nu,t}(\lambda))(\eta_{i}-\gamma)},\quad{\text{for each $1\leq i\leq p$}},$ (2.31)
    and $y_{0}(\gamma)=1-\sum_{i=1}^{p}y_{i}(\gamma)$. The local minimizers of $J^{\Lambda}_{\nu,t}(\lambda,.)$ are included in $\{Y(\omega_{\nu,t}^{*}(\lambda)),Y(\Lambda)\}$ whenever $y_{0}(\Lambda)\geq 0$ or equal to $\{Y(\omega_{\nu,t}^{*}(\lambda))\}$ whenever $y_{0}(\Lambda)<0$. Moreover, for $\gamma=\omega_{\nu,t}^{*}(\lambda)$ and $\gamma=\Lambda$, we have
@@
    The function $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ admits a unique minimizer $(y_{0},Y)\in\Delta^{p+1}$ in the following cases:
@@
    • If $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\in(\rho_{\nu,t}^{\Lambda},\ell_{\nu,t}]$, then the unique minimizer is $Y({\omega_{\nu,t}^{*}(\lambda)})$ with $y_{0}(\omega_{\nu,t}^{*}(\lambda))$ vanishing.
@@
Step 1: existence of a minimizer. By Cauchy-Schwarz inequality, since $t\geq 0$,
$J_{\nu,t}^{\Lambda}(\lambda,Y)\geq-\frac{\lambda}{2t}\Bigr{(}\Lambda y_{0}+\sum_{k=1}^{p}\eta_{k}y_{k}\Bigr{)}+\frac{1}{4t}\Bigr{(}\sum_{k=1}^{p}\eta_{k}^{2}y_{k}+\Lambda^{2}y_{0}\Bigr{)}-\frac{1}{2}\sum_{k=1}^{p}\alpha_{k}\log(y_{k}/\alpha_{k})$
@@
${J_{\nu,t}^{\Lambda}(\lambda,Y)}\geq\frac{1}{4t}\Bigr{(}\sum_{k=1}^{p}(\eta_{k}-\lambda)^{2}y_{k}+(\Lambda-\lambda)^{2}y_{0}\Bigr{)}-\frac{1}{4t}\lambda^{2}\geq-\frac{1}{4t}\lambda^{2},$
is uniformly bounded from below on $\Delta^{p+1}$. It is well-known that $Y\in(1-y_{0})\Delta^{p}\mapsto I_{\nu}(Y)$ is lower semicontinuous. Moreover,
@@
are continuous. Thus, $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ is lower semicontinuous and since it is bounded from below, it achieves its minimal value. The function $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ can be differentiated and its critical points can be described.
@@
Therefore, for each $k=1,\ldots,p$,
$y_{k}=\frac{t}{\eta_{k}^{2}-\lambda\eta_{k}-R_{1}\eta_{k}-R_{2}},$
@@
$y_{k}=\frac{t}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}{\quad\text{for each$k=1,\ldots,p$}}$ (2.34)
@@
It follows that $\gamma_{1}=\Lambda$ or $\gamma_{2}=\Lambda$ and $\gamma_{1},\gamma_{2}\in\mathbb{R}$. Thus, for each $k=1,\ldots,p$,
$y_{k}=\frac{t}{(\eta_{k}-\Lambda)(\eta_{k}-\alpha)},$
for some $\alpha\in\mathbb{R}$. Note that $y_{k}$ is nonnegative for every $k$ such that $\alpha$ belongs to $(-\infty,\ell_{\nu})$ since $\Lambda\leq\ell_{\nu}$. If $\alpha\neq\Lambda$, the first equation of (2.33) gives
$\sum_{k=1}^{p}y_{k}=\frac{t}{\Lambda-\alpha}(G_{\nu}(\alpha)-G_{\nu}(\Lambda))=1-y_{0},$
otherwise if $\alpha=\Lambda$, we have
@@
$H_{\nu,t}(\alpha)=\lambda,\quad{\alpha<\ell_{\nu}}.$ (2.35)
By Remark A.2, whenever $\lambda<\ell_{\nu,t}$, the above equation has two distinct solutions $\gamma_{1}<\gamma_{2}$ in $(-\infty,\ell_{\nu})$ and moreover $\gamma_{1}=\omega_{\nu,t}(\lambda)$ and $\gamma_{2}=\omega^{*}_{\nu,t}(\lambda)$. If $\lambda=\ell_{\nu,t}$, then Equation (2.35) has a unique solution $\alpha=\omega_{\nu,t}(\ell_{\nu,t})$.
@@
${y_{k}(\gamma)}:=\frac{t}{(\eta_{k}-\gamma)(\eta_{k}-\Lambda)}{\quad\text{for each $k=1,\ldots,p$}}.$ (2.36)
@@
which is strictly positive if and only if $\Lambda<\omega_{\nu,t}(\ell_{\nu,t})$. Therefore if $\lambda=\rho_{\nu,t}^{\Lambda}=H_{\nu,t}(\Lambda)$ (giving $\gamma_{1}=\gamma_{2}=\Lambda$) and $\Lambda<\omega_{\nu,t}(\ell_{\nu,t})$, then $y_{0}\in(0,1)$. If ${\Lambda\notin\{\gamma_{1},\gamma_{2}\}}$, then
@@
and $y_{0}({\gamma})$ belongs to $(0,1)$ only if
@@
Hence, in this range of parameters $Y({\gamma_{i}})_{i=1,2}$ are critical points in $\Delta^{p+1}$, but none of them belong to $\Delta^{p+1}$ otherwise. In the following we compute the value of $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ for each of these two solutions in order to select the minimizer. Let $\gamma\in\{\gamma_{1},\gamma_{2}\}$. One may compute
$X:=\Lambda y_{0}+\sum_{i=1}^{p}\eta_{i}y_{i}=\Lambda-tG_{\mu}(\gamma)=\Lambda+\gamma-\lambda,$ (2.38)
@@
$\begin{split}\Lambda^{2}y_{0}+\sum_{k=1}^{p}\eta_{k}^{2}y_{k}&=\Lambda^{2}y_{0}+t+(\gamma+\Lambda)\sum\eta_{k}y_{k}-\gamma\Lambda(1-y_{0})\\ &=\Lambda^{2}y_{0}+t+(\gamma+\Lambda)X-\gamma\Lambda(1-y_{0})-(\gamma+\Lambda)\Lambda y_{0}\\ &=t+(\gamma+\Lambda)X-\gamma\Lambda.\end{split}$
Thus
$\begin{split}J_{\nu,t}^{\Lambda}(\lambda,Y(\gamma))&=\frac{1}{4t}(-2\lambda-X+2(\gamma+\Lambda))X+\frac{1}{2}-\frac{\gamma\Lambda}{2t}-\frac{1}{2}(S_{\nu}(\Lambda)+S_{\nu}(\gamma))-\frac{1}{2}\ln t\\ &=\frac{1}{4t}(\gamma+\Lambda-\lambda)^{2}-2\gamma\Lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\Lambda)+S_{\nu}(\gamma))-\frac{1}{2}\ln t\\ &=\frac{1}{4t}(\gamma^{2}+\Lambda^{2}+\lambda^{2}-2\gamma\lambda-2\Lambda\lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\Lambda)+S_{\nu}(\gamma))-\frac{1}{2}\ln t.\end{split}$ (2.39)
@@
Step 3: study of the critical points on the boundary. Let $Y\in\Delta^{p}$ be a critical point of $J_{\nu,t}^{\Lambda}(\cdot,0,Y)$. By variational calculus, there exist $R_{1},R_{2}$ such that
$-\lambda\eta_{k}+\eta_{k}^{2}-R_{1}\eta_{k}-\frac{t}{y_{k}}=R_{2},{\quad\text{for each $1\leq k\leq p$}},$
@@
$y_{k}=\frac{t}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})},{\quad\text{for each $1\leq k\leq p$}}$ (2.40)
@@
Therefore $\gamma_{1}=\omega_{\nu,t}(\lambda)$ and $\gamma_{2}=\omega_{\nu,t}^{*}(\lambda)$. Let us now compute $J_{\nu,t}^{\Lambda}(\lambda,0,Y)$. Let $X:=\sum\eta_{k}y_{k}=\gamma_{1}+\gamma_{2}-\lambda$. We have
${\sum_{k=1}^{p}}\eta_{k}^{2}y_{k}=t+(\gamma_{1}+\gamma_{2})X-\gamma_{1}\gamma_{2}.$
@@
$\begin{split}J_{\nu,t}^{\Lambda}(\lambda,0,Y)&=\frac{1}{4t}(-2\lambda-X+2(\gamma_{1}+\gamma_{2}))X+\frac{1}{2}-\frac{\gamma_{1}\gamma_{2}}{2t}-\frac{1}{2}(S_{\nu}(\gamma_{1})+S_{\nu}(\gamma_{2}))-\frac{1}{2}\ln t\\ &=\frac{1}{4t}(\gamma_{1}^{2}+\gamma_{2}^{2}+\lambda^{2}-2\gamma_{1}\lambda-2\gamma_{2}\lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\gamma_{1})+S_{\nu}(\gamma_{2}))-\frac{1}{2}\ln t.\end{split}$ (2.42)
Step 4: Identification of the minimizers.
@@
    $y_{k}(\Lambda)=\frac{t}{(\eta_{k}-\omega_{\nu,t}(\lambda))(\eta_{k}-\Lambda)},\quad{\text{for each $k=1,\ldots,p$}}.$
@@
By Lemma 2.4 and the proof of the previous lemma, it is enough to prove that $\Phi(\Lambda,Y(\Lambda))=\omega_{\nu,t}^{*}(\lambda)$ where we recall that $\Phi(\Lambda,Y)$ is defined as the smallest solution $\Lambda_{1}$ in (2.12). We denote in short $\alpha=\omega_{\nu,t}(\lambda)$. Since $y_{0}(\Lambda)\neq 0$, we may assume $\Lambda_{1}\neq\Lambda$. We claim that also $\Lambda_{1}\neq\alpha$ because
@@
$\begin{split}\sum_{k=0}^{p}\frac{y_{k}}{\eta_{k}-\alpha}&=\frac{y_{0}}{\Lambda-\alpha}-\frac{t}{(\Lambda-\alpha)^{2}}(-G_{\mu}(\alpha)+G_{\mu}(\Lambda)-(\Lambda-\alpha)G_{\mu}^{\prime}(\alpha))+O(\varepsilon_{n})\\ &=-\frac{1}{(\Lambda-\alpha)^{2}}(H_{\nu,t}(\alpha)-H_{\nu,t}(\Lambda)-tG_{\mu}(\alpha)+tG_{\mu}(\Lambda)-t(\Lambda-\alpha)G_{\mu}^{\prime}(\alpha))+O(\varepsilon_{n})\\ &=\frac{1+tG_{\nu}^{\prime}(\alpha)}{\Lambda-\alpha}.\end{split}$

### v2

@@ 2.6. Solving the functional equation in the complement of the region $\Lambda<\omega_{\nu,t}(\lambda)$
where $\omega^{*}_{\nu,t}(\lambda)$ is as in (1.15) and $\rho_{\nu,t}^{\Lambda}$ as in (1.13).Also define
@@
and recall ${I_{\nu,t}^{\Lambda}(\lambda):=I_{\nu,t}^{\Lambda}(\lambda,\gamma(\Lambda,\lambda))}$. We first prove that $F_{\nu,t}^{+}$ and $F_{\nu,t}^{-}$ equal $I_{\nu,t}^{\Lambda}$ in the simplest case where the functional equations (2.22) and (2.33) are such that the infimum is taken at $Y$ so that the term depending on $F_{\nu,t}^{+}$ or $F_{\nu,t}^{-}$ vanishes.
@@
Let $\rho_{\nu,t}^{\Lambda}$ be as in (1.13). We have for every $\lambda<\ell_{\nu,t}$ and $\Lambda\leq\ell_{\nu}$,
@@
Moreover, if ${\Lambda\in[\omega_{\nu,t}(\ell_{\nu,t}),\ell_{\nu}]}$ or $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$, then
@@
(a) $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$
(b) $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$
(c) $\Lambda<\omega_{\nu,t}(\lambda)$
Figure 2. The graph of the function $x\in(-\infty,\ell_{\nu})\mapsto x+tG_{\nu}(x)$.
Notice that the value of $I_{\nu,t}^{\Lambda}$ is computed in cases $A$ and $B$ of Figure 2. The case that remains to study is case $C$ where $\Lambda<\omega_{\nu,t}(\ell_{\nu,t})$. A key point to prove Lemma 2.8 is to study the minimizers of $K_{\nu,t}(\Lambda,\lambda,.)$, namely the minimizers of the function $J_{\nu,t}^{\Lambda}(\lambda,.)$ defined in (2.20). We show that in case $A$ and $B$, the minimizer $Y$ of $J_{\nu,t}^{\Lambda}(\lambda,.)$ is known, which crucially uses the fact that $\Phi(\Lambda,Y)\geq\omega_{\nu,t}(\ell_{\nu,t})$, where $\Phi(\Lambda,Y)$ is as in (2.14). We will prove Lemma 2.8 after this identification of the minimizer of $J_{\nu,t}^{\Lambda}(\lambda,.)$.
@@
Fix $\lambda\in(-\infty,\ell_{\nu,t}]$. Let $\rho_{\nu,t}^{\Lambda}$ be as in (1.13) and $\omega_{\nu,t}^{*}$ be as in (1.15).
@@
    $y_{i}(\gamma):=\frac{t{\alpha_{i}}}{(\eta_{i}-\omega_{\nu,t}(\lambda))(\eta_{i}-\gamma)},\quad{\text{for {every} $1\leq i\leq p$}},$ (2.39)
    and $y_{0}(\gamma)=1-\sum_{i=1}^{p}y_{i}(\gamma)$. The local minimizers of $J^{\Lambda}_{\nu,t}(\lambda,.)$ are included in $\{Y(\omega_{\nu,t}^{*}(\lambda)),Y(\Lambda)\}$ whenever $y_{0}(\Lambda)\geq 0$ or equal to $Y(\omega_{\nu,t}^{*}(\lambda))=\{(0,(y_{i}(\omega_{\nu,t}^{*}(\lambda)))_{1\leq i\leq p})\}$ whenever $y_{0}(\Lambda)<0$. Moreover, for $\gamma=\omega_{\nu,t}^{*}(\lambda)$ and $\gamma=\Lambda$, we have
@@
    The function $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ admits a unique minimizer $Y\in\Delta^{p+1}$ in the following cases:
@@
    • If $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\in(\rho_{\nu,t}^{\Lambda},\ell_{\nu,t}]$, then the unique minimizer is $Y({\omega_{\nu,t}^{*}(\lambda)})$ with $y_{0}(\omega_{\nu,t}^{*}(\lambda))$ equal to zero.
@@
Step 1: existence of a minimizer. By Cauchy-Schwarz inequality, since $t\geq 0$ and $Y\in\Delta^{p+1}$,
$J_{\nu,t}^{\Lambda}(\lambda,Y)\geq-\frac{\lambda}{2t}\Bigr{(}\Lambda y_{0}+\sum_{k=1}^{p}\eta_{k}y_{k}\Bigr{)}+\frac{1}{4t}\Bigr{(}\sum_{k=1}^{p}\eta_{k}^{2}y_{k}+\Lambda^{2}y_{0}\Bigr{)}-\frac{1}{2}\sum_{k=1}^{p}\alpha_{k}\log(y_{k}/\alpha_{k})\,.$
@@
${J_{\nu,t}^{\Lambda}(\lambda,Y)}\geq\frac{1}{4t}\Bigr{(}\sum_{k=1}^{p}(\eta_{k}-\lambda)^{2}y_{k}+(\Lambda-\lambda)^{2}y_{0}\Bigr{)}-\frac{1}{4t}\lambda^{2}\geq-\frac{1}{4t}\lambda^{2}.$
Therefore $J_{\nu,\Lambda}(\lambda,\cdot)$ is uniformly bounded from below on $\Delta^{p+1}$. It is well-known that $Y\in(1-y_{0})\Delta^{p}\mapsto I_{\nu}(Y)$ is lower semicontinuous. Moreover,
@@
are continuous. Thus, $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ is lower semicontinuous and since it is bounded from below, it achieves its minimal value. The function $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ can be differentiated and its critical points analyzed. We describe them in the sequel.
@@
Therefore, for every $k=1,\ldots,p$,
$y_{k}=\frac{t\alpha_{k}}{\eta_{k}^{2}-\lambda\eta_{k}-R_{1}\eta_{k}-R_{2}},$
@@
$y_{k}=\frac{t\alpha_{k}}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}{\quad\text{for {every} $k=1,\ldots,p$}}$ (2.42)
@@
It follows that $\gamma_{1}=\Lambda$ or $\gamma_{2}=\Lambda$ and $\gamma_{1},\gamma_{2}\in\mathbb{R}$. Thus, for every $k=1,\ldots,p$,
$y_{k}=\frac{t\alpha_{k}}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma)},$
for some $\gamma\in\mathbb{R}$. Note that $y_{k}$ is nonnegative for every $k$ such that $\gamma$ belongs to $(-\infty,\ell_{\nu})$ since $\Lambda\leq\ell_{\nu}$. If $\gamma\neq\Lambda$, the first equation of (2.41) gives
$\sum_{k=1}^{p}y_{k}=\frac{t}{\Lambda-\gamma}(G_{\nu}(\gamma)-G_{\nu}(\Lambda))=1-y_{0},$ (2.43)
otherwise if $\gamma=\Lambda$, we have
@@
$H_{\nu,t}(\gamma)=\lambda,\quad{\gamma<\ell_{\nu}}.$ (2.44)
By Remark A.2, whenever $\lambda<\ell_{\nu,t}$, the above equation has two distinct solutions $\gamma_{1}<\gamma_{2}$ in $(-\infty,\ell_{\nu})$ and moreover $\gamma_{1}=\omega_{\nu,t}(\lambda)$ and $\gamma_{2}=\omega^{*}_{\nu,t}(\lambda)$. If $\lambda=\ell_{\nu,t}$, then Equation (2.44) has a unique solution $\gamma=\omega_{\nu,t}(\ell_{\nu,t})$.
@@
${y_{k}(\gamma)}:=\frac{t\alpha_{k}}{(\eta_{k}-\gamma)(\eta_{k}-\Lambda)}{\quad\text{{for every} $k=1,\ldots,p$}}.$ (2.45)
@@
which is strictly positive if and only if $\Lambda<\omega_{\nu,t}(\ell_{\nu,t})$. Therefore if $\lambda=\rho_{\nu,t}^{\Lambda}=H_{\nu,t}(\Lambda)$ (giving $\gamma_{1}=\gamma_{2}=\Lambda$) and $\Lambda<\omega_{\nu,t}(\ell_{\nu,t})$, then $y_{0}\in(0,1)$. If ${\Lambda\notin\{\gamma_{1},\gamma_{2}\}}$, then (2.46) shows that
@@
and $y_{0}({\gamma})$ is positive only if
@@
Note that $y_{0}(\gamma)\leq 1$ by (2.43) since $\alpha$ and $\Lambda$ are smaller or equal to $\ell_{\nu}$. Hence, in this range of parameters $Y({\gamma_{i}})_{i=1,2}$ are critical points in $\Delta^{p+1}$, but none of them belong to $\Delta^{p+1}$ otherwise. In the following we compute the value of $J_{\nu,t}^{\Lambda}(\lambda,\cdot)$ for every of these two solutions in order to select the minimizer. Let $\gamma\in\{\gamma_{1},\gamma_{2}\}$. One may compute
$M(\gamma):=\Lambda y_{0}(\gamma)+\sum_{i=1}^{p}\eta_{i}y_{i}(\gamma)=\Lambda-tG_{\mu}(\gamma)=\Lambda+\gamma-\lambda,$ (2.47)
@@
$\begin{split}\sigma(\gamma)&:=\Lambda^{2}y_{0}(\gamma)+\sum_{k=1}^{p}\eta_{k}^{2}y_{k}(\gamma)\\ &=\Lambda^{2}y_{0}(\gamma)+t\sum_{k=1}^{p}\frac{\alpha_{k}}{(\eta_{k}-\gamma)(\eta_{k}-\Lambda)}\left((\eta_{k}-\gamma)(\eta_{k}-\Lambda)+(\gamma+\Lambda)\eta_{k}-\gamma\Lambda\right)\\ &=\Lambda^{2}y_{0}(\gamma)+t+(\gamma+\Lambda)\sum\eta_{k}y_{k}(\gamma)-\gamma\Lambda(1-y_{0}(\gamma))\\ &=t+(\gamma+\Lambda)M(\gamma)-\gamma\Lambda.\end{split}$
Moreover, since $\sum\alpha_{k}=1$,
$I_{\nu}(Y(\gamma))=-\frac{1}{2}\sum_{k=1}^{p}\alpha_{k}\ln\left(\frac{t}{(\eta_{k}-\gamma)(\eta_{k}-\Lambda)}\right)=-\frac{1}{2}\left(S_{\nu}(\gamma)+S_{\nu}(\Lambda)\right)-\frac{1}{2}\ln t\,.$ (2.48)
Thus, we find
$\begin{split}J_{\nu,t}^{\Lambda}(\lambda,Y(\gamma))&=L^{\Lambda}_{\nu,t}(Y(\gamma))+I_{\nu}(Y(\gamma))=-\frac{\lambda}{2t}M(\gamma)+\frac{1}{2t}\sigma(\gamma)-\frac{1}{4t}M(\gamma)^{2}+I_{\nu}(Y(\gamma))\\ &=\frac{1}{4t}(-2\lambda-M(\gamma)+2(\gamma+\Lambda))M(\gamma)+\frac{1}{2}-\frac{\gamma\Lambda}{2t}-\frac{1}{2}(S_{\nu}(\Lambda)+S_{\nu}(\gamma))-\frac{1}{2}\ln t\\ &=\frac{1}{4t}(\gamma+\Lambda-\lambda)^{2}-2\gamma\Lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\Lambda)+S_{\nu}(\gamma))-\frac{1}{2}\ln t\\ &=\frac{1}{4t}(\gamma^{2}+\Lambda^{2}+\lambda^{2}-2\gamma\lambda-2\Lambda\lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\Lambda)+S_{\nu}(\gamma))-\frac{1}{2}\ln t.\end{split}$ (2.49)
@@
Step 3: study of the critical points on the boundary. We next consider the critical points so that $y_{0}$ equal to zero. Let $Y\in\Delta^{p}$ be a critical point of ${J_{\nu,t}^{\Lambda}(\lambda,0,\cdot)}$. By variational calculus, there exist $R_{1},R_{2}$ such that
$-\lambda\eta_{k}+\eta_{k}^{2}-R_{1}\eta_{k}-\frac{t\alpha_{k}}{y_{k}}=R_{2},{\quad\text{{for every} $1\leq k\leq p$}},$
@@
$y_{k}=\frac{t\alpha_{k}}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})},{\quad\text{{for every} $1\leq k\leq p$}}$ (2.50)
@@
Therefore $\gamma_{1}=\omega_{\nu,t}(\lambda)$ and $\gamma_{2}=\omega_{\nu,t}^{*}(\lambda)$ where $\omega_{\nu,t}^{*}(\lambda)$ is as in (1.15). Let us now compute $J_{\nu,t}^{\Lambda}(\lambda,0,Y)$. Let $M(\gamma):=\sum\eta_{k}y_{k}(\gamma)=\gamma_{1}+\gamma_{2}-\lambda$. We have
${\sum_{k=1}^{p}}\eta_{k}^{2}y_{k}=t+(\gamma_{1}+\gamma_{2})M(\gamma)-\gamma_{1}\gamma_{2}.$
@@
$\begin{split}J_{\nu,t}^{\Lambda}(\lambda,0,Y)&=\frac{1}{4t}(-2\lambda-M(\gamma)+2(\gamma_{1}+\gamma_{2}))M(\gamma)+\frac{1}{2}-\frac{\gamma_{1}\gamma_{2}}{2t}-\frac{1}{2}(S_{\nu}(\gamma_{1})+S_{\nu}(\gamma_{2}))-\frac{1}{2}\ln t\\ &=\frac{1}{4t}(\gamma_{1}^{2}+\gamma_{2}^{2}+\lambda^{2}-2\gamma_{1}\lambda-2\gamma_{2}\lambda)+\frac{1}{2}-\frac{1}{2}(S_{\nu}(\gamma_{1})+S_{\nu}(\gamma_{2}))-\frac{1}{2}\ln t.\end{split}$ (2.52)
Step 4: Identification of the minimizer.
@@
    $y_{k}(\Lambda)=\frac{t\alpha_{k}}{(\eta_{k}-\omega_{\nu,t}(\lambda))(\eta_{k}-\Lambda)},\quad{\text{{for every} $k=1,\ldots,p$}}.$
@@
Let us prove (1). By Lemma 2.4 and the proof of the previous lemma, it is enough to prove that $\Phi(\Lambda,Y(\Lambda))=\omega_{\nu,t}^{*}(\lambda)$ where we recall that $\Phi(\Lambda,Y)$ is defined as the smallest solution $\Lambda_{1}$ in (2.15). We denote in short $\alpha=\omega_{\nu,t}(\lambda)$. Since $y_{0}(\Lambda)\neq 0$, we may assume $\Lambda_{1}\neq\Lambda$. We claim that also $\Lambda_{1}\neq\alpha$ because
@@
$\begin{split}\sum_{k=0}^{p}\frac{y_{k}}{\eta_{k}-\alpha}&=\frac{y_{0}}{\Lambda-\alpha}-\frac{t}{(\Lambda-\alpha)^{2}}(-G_{\mu}(\alpha)+G_{\mu}(\Lambda)-(\Lambda-\alpha)G_{\mu}^{\prime}(\alpha))\\ &=-\frac{1}{(\Lambda-\alpha)^{2}}(H_{\nu,t}(\alpha)-H_{\nu,t}(\Lambda)-tG_{\mu}(\alpha)+tG_{\mu}(\Lambda)-t(\Lambda-\alpha)G_{\mu}^{\prime}(\alpha))\\ &=\frac{1+tG_{\nu}^{\prime}(\alpha)}{\Lambda-\alpha}.\end{split}$